there are 25 coins in a box, some dimes, some quarters. They have a value of $3.70. How many of each coin?
Answer by mananth(16949)

let dimes be x
quarters will be 25-x
..
total value =$3.70 = 370 cents
...
10x+25(25-x)= 370
10x+625-25x=370
-15x=-625+370
-15x=-255
/-15
-15x/-15 = -255/-15
x= 17 ----- dimes
quarters will be 25-17 = 8
...
m.ananth@hotmail.ca